OpenStudy (anonymous):

Please Help A tray contains 14 chocolate chip cookies and 16 oatmeal raisin cookies. How many ways can a child select 4 cookies that include at least 1 chocolate chip cookie?

OpenStudy (anonymous):

Can someone help me please?

OpenStudy (anonymous):

at least one means one or two or three or four

OpenStudy (anonymous):

Try using a binomial distribution.

OpenStudy (anonymous):

total - no chocolate chip cookies

OpenStudy (anonymous):

(30C4) - (14C4)

OpenStudy (anonymous):

4 cookies lets say all are chocolate chip the number of ways to choose 4 out of 14 is written \(\binom{14}{4}\) or sometimes \(_{14}C_4\) or even \(^{16}C_4\) and computed via \[\binom{14}{4}=\frac{\overbrace{14\times 13\times 12\times 11}^{\text{ 4 terms}}}{4\times 3\times 2}\] cancel first , multiply last since it is a whole number

OpenStudy (anonymous):

actually @sourwing method is better, but i think there is a typo there

OpenStudy (anonymous):

oh yeah I meant (30C4) - (16C4) :DD